OpenStudy (anonymous):

Which value of x satisfies both -9x + 4y = 8 and -3x − y = 4?

OpenStudy (anonymous):

A. -1/7 B. -7/9 C. -8/7 D. -9 E. -4/3

OpenStudy (aravindg):

Plugin and see!

OpenStudy (anonymous):

Whats the answer?

OpenStudy (anonymous):

Multiply the second equation by 4: 4 * (-3x - y) = 4 * 4 -12x - 4y = 16 Now add the two equations together: -9x + 4y = 8 + -12x - 4y = 16 -21x + 0y = 24 x = 24/-21 x = -8/7 The answer is C.